OpenStudy (anonymous):

Factor completely 9x^2 + 9x -28

myininaya (myininaya):

9(-28)=21*(-12) and 21-12=9 so we can factor this poly. 9x=21x-12x we will use this substitution ^ | 9x^2+21x-12x-28 =3x(3x+7)-4(3x+7) =(3x+7)(3x-4)